OpenStudy (anonymous):

Medal A rifle bullet travels 5000 feet in 4 seconds .What is the speed Of the bullet ?

OpenStudy (anonymous):

V=D/T velocity equals distance divided by time. So V=6000/4. V=1500 I hope this is right!